From GlueXWiki
Jump to: navigation, search

August 27, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Electronics
  2. CDC update (Mike, Simon, Beni)
  3. FDC update (Luke,Lubomir)
  4. Engineering
  5. Other


Participants: Curtis, Mike, Naomi (CMU), Fernando, Eugene, Luke, Beni, Dave, Chris, Nick, Simon, Sergey and Lubomir (JLab).


- Fernando studied the noise on the CDC and FDC together using two differential probes: see scope pictures linked in the agenda. First picture shows the noise levels in the two detectors are similar; FDC looks better but not by a lot. The second and third pictures show typical cosmic signals with correct timing. Last plot discussions - occasionally there are long tails in the CDC signals (~5 usec). The signals are still low enough (~100mV) so this is not a saturation effect (Fernando). Most likely it comes from ion tails (Lubomir), however it is not clear why it happens mostly in the CDC, is the ion-cancellation different in the two detectors?

- Nick: all "noise-reducing" elements from the CDC were removed. Tim is informed about that but needs to decide about the installation of an additional Al-foil target before inserting the target back. Nick replaced one LV module in the CDC; it was giving high current without any load. This is the second problematic LV module, the first one had different issue.

- The PR for 100 spare pre-amps was approved.

CDC update

- Simon continues to study the tube deformations using beam data. So far he checked rings 1 to 7 and 13. Ring 7 (stereo) looks worse and there's a straw with more than 1mm deformation.

- Beni informed us that the Bryan's libraries on the hdcdcops account are the new ones, corresponding to the new f125 firmware.

FDC update

- Lubomir showed the change of the spring 2015 alignment offsets w,r.t. fall 2014 run (plot attached - black/red are x/y offsets for each cell given by the z-coordinate). The max deviations are for the first package +/- 200 microns. Most likely it has to do with the alignment procedure, that starts with package 3 then 4-2-1. The problem is however, that with the new alignments the resolutions in the spring runs are worse than in the fall run, like ~250 microns compared to ~200 microns in the fall. May have to do with the lower energies.

- Luke presented his latest results about the avalanche position studies with 55Fe. First plot shows the avalanche position in plane perpendicular to the wire - empty area around the wire is visible. Second plot: the two maximums in up/down direction vs gain - distance between them increases with decreasing the gain (no idea why!). Last plot is the same but for the maximums in left/right direction - no dependence on the gain as expected.


- We have now two new MFC calibrated for 2 l/min max flow. Scot will replace the bad CDC CO2 MFC and we will separate the two gas systems. Scot is also installing additional relief valves at the high pressure lines.